I am working on this problem and I think I know some of the steps. I'll show you what I have done so far. 1/2x + 1 = 3/4
1/2x +1 -1= 3/4 - 4/4
1/2x = -1/4
where do I go from here?

Answer by Alan3354(69443)

I am working on this problem and I think I know some of the steps. I'll show you what I have done so far. 1/2x + 1 = 3/4
1/2x +1 -1= 3/4 - 4/4
1/2x = -1/4
where do I go from here?
------------
If it's (1/2)x,
Multiply both side by 2
x = -1/2